1. A typical family on Sandy Island consumes only juice and cloth. Last year, which was the base year, the family spent $40 on juice and $25 on cloth. In the base year, juice $4 a bottle and cloth is $5 a length. This year, juice is $4 a bottle and cloth is $6 a length. Calculate:
  2. The CPI basket
  3. The CPI in the current year
  4. The inflation rate in the current year
1
Expert's answer
2022-02-24T11:44:23-0500

(a)CPI basket;

no of juices purchased="\frac {money spent on juice}{price of juice bottle}" ="\frac{40}{4}=10"

no of clothe purchased"=\frac{ no of cloth purchased}{price of clothe}=\frac {25}{5}=5"

so CPI basket is 10 bottle of juice and 5 units of cloth

(b). CPI of current year"={price of juice} \times{no of juice bottle}+{price of cloth} \times{no of cloth unit}=" 4"\times" 10+6"\times"5=40+30=70

(C).cpi in base year=money spent on juice + money spent on cloth=$40+$25=$65

.inflation rate = "(\frac{cpi in current year -cpi in base year}{cpi in base year})\times 100" ="(\frac {70-65}{65})\times100" =7.69

thus inflation rate =7.69%